# Dual MPC for Active Learning of Nonparametric Uncertainties

### v1

@@ Abstract
This manuscript presents a dual model predictive controller (DMPC) that balances the two objectives of dual control, namely, system identification and control. In particular, we propose a Gaussian process (GP)-based MPC that uses the posterior GP covariance for active learning. The DMPC can steer the system towards states that have high covariance, or to the setpoint, thereby balancing system identification and control performance (exploration vs. exploitation). We establish robust constraint satisfaction of the novel DMPC through the use of a contingency plan. We demonstrate the DMPC in a numerical study of a nonlinear system with nonparametric uncertainties.
@@ I Introduction
One of the reasons for the high interest in model predictive control (MPC), in particular for safety-critical systems, is its ability to systematically handle constraints on states and inputs. Yet, the performance and safety of MPC heavily depend on the accuracy of the model of the system. In various applications, the system dynamics may be uncertain or partially unknown, and need to be identified online. This requires sufficiently informative data, and hence, the excitation of the system. However, these system excitations may conflict with the control objective and safety requirements. Balancing these objectives, identification and control, is the focus of dual control [FELDBAUM1963541]. Therefore, studying dual control in the context of MPC is of high relevance for simultaneous identification and control of safety-critical systems.
@@ I-A The Dual Effect
When controlling a system that is subject to uncertainties, a controller can only make decisions based on the available information of the system. However, the possibility of causally anticipating future measurements can be leveraged to improve estimation or prediction, and hence, control performance. When a controller incorporates a ‘future measurement program’ and its associated statistics, this controller can causally anticipate these future measurements [bar-shalom_dual_1974]. The control is said to have the dual effect, when, in addition to influencing the system state, it can also influence the state uncertainty, e.g., the state covariance. As a consequence, the dual effect can be comprised of two properties [bar-shalom_dual_1974]: 1) Caution: the dependence of the state covariance on the control policy naturally introduces a degree of inherent caution – or a lack thereof. In the context of MPC, this dependence can be leveraged to tighten or relax the state constraints based on the anticipated state covariance. 2) Exploration: this anticipated state covariance can also be used to introduce an explorative or exciting element to the control policy. Again, in the context of MPC, the state covariance can be used in the MPC cost function to steer towards regions of high uncertainty that may be of interest for system identification and control performance. For a given control problem, the solution to the Bellman equation defines the globally optimal control policy, which naturally balances identification and control in an optimal way. In practice, however, solving the Bellman equation is generally intractable and requires approximate solutions [mesbah_stochastic_2018].
@@ I-B1 Dual Control
In implicit dual control, the MPC cost function is used to approximate the Bellman equation. These types of methods are typically associated with and related to approximate dynamic programming. The trade-off between exploration and exploitation is implicitly governed by the cost function. For example, [arcari_dual_2020a] and [HUSHARP2024] use scenario trees in MPC to address the exploration-exploitation trade-off through dynamic programming. However, existing methods rely on additional safety filters [HUSHARP2024], are limited to linear systems [soloperto_dual_2019], or do not consider state constraints [arcari_dual_2020a], which is essential for safety-critical applications. The implementation and computational complexity of implicit dual MPC limit its applicability [mesbah_stochastic_2018].
In explicit dual control, the explorative element of the control policy is introduced explicitly, e.g., by enforcing a persistently exciting input signal, or adding a term to the cost function that minimizes (future) model uncertainty [mesbah_stochastic_2018]. We refer to [mesbah_stochastic_2018] for a detailed review of these methods. Although the exploration in explicit dual MPC typically requires some form of expert knowledge and is often heuristic [mesbah_stochastic_2018], these methods can explicitly account for state constraints. Therefore, explicit dual MPC is of high interest, especially for safety-critical applications. While existing methods are highly relevant, they mostly focus on systems with parametric uncertainties [mesbah_stochastic_2018] or fixed model structures [heirung_stochastic_2017]. This raises the question of how to effectively extend these methods to address nonparametric model uncertainties in dual MPC.
@@ I-B2 Gaussian Process MPC
Gaussian process (GP) regression is a Bayesian regression method that provides a systematic way for inductive reasoning and nonparametric modeling that can be leveraged with limited expert knowledge and few data points [kocijan_modelling_2016]. It has been successfully used to identify and control systems with unknown dynamics using MPC [hewing_cautious_2020, baltussen2025, dubied2025robustadaptivempcformulation]. While GP-MPC is often applied as an effective heuristic method [hewing_cautious_2020, baltussen2025], these methods can be complemented with robustly safe policies [geurts_baltussen_2025] that ensure constraint satisfaction under uncertainty. Alternatively, probabilisitic safety guarantees on GP-MPC are provided in [dubied2025robustadaptivempcformulation]. However, existing GP-MPC methods generally lack the dual effect as the GP covariance is not conditioned on the predicted control sequence. Typically the GP covariance is fixed prior to solving the MPC problem [hewing_cautious_2020] or decoupled from the MPC problem in a zero-order optimization algorithm [dubied2025robustadaptivempcformulation] to enable real-time implementation of GP-MPC. While of great interest, decoupling the GP covariance prevents the dual control effect. To bridge this gap, [baltussen2025] introduced a GP-MPC that conditions the GP covariance on the predicted control sequence, yielding a policy with the dual effect, which is a neccessary property for active learning with GP-MPC.
@@ I-C Contributions
As mentioned before, in [baltussen2025] we introduced a GP-MPC that features the dual effect. However, we did not explicitly incentivize exploration. Hence, this passive learning dual MPC was limited to the cautious property of the dual effect. In this work, we extend the GP-MPC to actively explore the state space and to learn unknown system dynamics while ensuring robust constraint satisfaction. The main contributions of this manuscript are as follows. 1) We propose an explicit dual MPC that actively learns system dynamics using the active learning framework from [soloperto_augmenting_2020]. 2) We prove recursive feasibility of the active learning MPC by incorporating a contingency horizon[alsterda_contingency_2019]. 3) We demonstrate the effectiveness of the proposed dual MPC through a numerical study on a nonlinear system with nonparametric, unknown dynamics. In addition, we compare our dual MPC with a robust MPC (RMPC), a passively learning dual MPC [geurts_baltussen_2025], and the active learning framework from [soloperto_augmenting_2020] with a single horizon.
@@ III-A Control Policies and the Dual Effect
According to Definition 2, only the $N$-measurement feedback and closed-loop control policies exhibit this dual effect [bar-shalom_dual_1974, mesbah_stochastic_2018]. This effect is a necessary property for active learning in GP-MPC, as it allows the MPC to actively influence the GP posterior covariance and guide the system toward uncertain regions to learn the dynamics and enhance performance.
@@ III-B Robust Model Predictive Control
$\bar{u}_{i\mid k}$ $\in\bar{\mathbb{U}}_{i},\ \ \ \ \;i=0,1,\dots,N-1,$ (3d)
$\bar{x}_{i\mid k}$ $\in\bar{\mathbb{X}}_{i},\ \ \ \ \;i=0,1,\dots,N,$ (3e)
@@ III-C1 Gaussian Process Regression
$y_{k+1}=B^{{\dagger}}\left(x_{k+1}-f(x_{k},u_{k})\right)=g\left(x_{k},u_{k}\right)+\tilde{w}_{k}$
with $B^{{\dagger}}\hskip-4.2679pt=\hskip-1.42262pt\left(B^{\top}B\right)^{-1}B^{\top}$ and $B^{{\dagger}}v_{k}=\tilde{w}_{k}\sim\mathcal{N}(0,\Sigma_{w})$. We use the training data $\mathbf{Z}_{k}=[\mathbf{z}_{0},\dots,\mathbf{z}_{k-1}]\in\mathbb{R}^{n_{z}\times k}$ and $\mathbf{y}_{k}=[y_{1},\dots,y_{k}]\in\mathbb{R}^{n_{g}\times k}$} at time $k\in\mathbb{N}$ in order to approximate the residual dynamics by a GP $\tilde{\mathbf{d}}\hskip-1.42262pt:\mathbb{R}^{n_{z}}\rightarrow\mathbb{R}^{n_{g}}$ such that $g\left(x_{k},u_{k}\right)\approx\tilde{\mathbf{d}}\left(\mathbf{z}_{k}\right)$. We confine ourselves to scalar GPs $(n_{g}=1)$ for the sake of readability. Secondly, we impose a prior distribution on the GP through a universal kernel function $\mathfrak{K}\left(\mathbf{z},\mathbf{z}^{\prime}\right)$ [kocijan_modelling_2016]. As we use the GP for model augmentation, we employ a zero mean prior. Conditioning the prior GP on the information $\mathcal{I}_{k}$ yields the posterior GP. Finally, we predict the residual dynamics $g$ over a horizon $i=0,1,\dots,N-1$ at time $k\in\mathbb{N}$ by querying the GP posterior $\tilde{d}(\mathbf{z})\in\mathbb{R}$, at a predicted test point $\mathbf{z}_{i\mid k}\in\mathbb{R}^{n_{z}}$ with
@@
$d(\mathbf{z})$ $=\mathbf{k}_{\mathbf{zZ}}\left(K_{\mathbf{ZZ}}+I\sigma^{2}_{w}\right)^{-1}\mathbf{y},$ (5a)
$\Sigma^{d}(\mathbf{z})$ $=\mathfrak{K}_{\mathbf{zz}}-\mathbf{k}_{\mathbf{zZ}}\left(K_{\mathbf{ZZ}}+I\sigma^{2}_{w}\right)^{-1}\mathbf{k}_{\mathbf{Z}\mathbf{z}}.$ (5b)
Here $\mathfrak{K}_{\mathbf{zz}^{\prime}}=\mathfrak{K}\left(\mathbf{z},\mathbf{z}^{\prime}\right)$, $\mathbf{k}_{\mathbf{Zz}}\in\mathbb{R}^{{k}}$ is the concatenation of the kernel function evaluated at the test point $\mathbf{z}$ and the training set $\mathbf{Z}$, where $\mathbf{k}_{\mathbf{Zz}}]_{i}=\mathfrak{K}\left(\mathbf{z}_{i},\mathbf{z}\right)$ and $\mathbf{k}_{\mathbf{Zz}}^{\top}=\mathbf{k}_{\mathbf{zZ}}$}, and $K_{\mathbf{ZZ}}\in\mathbb{R}^{{k}\times{k}}$ is a Gram matrix, which satisfies $K_{\mathbf{ZZ^{\prime}}}]_{ij}=\mathfrak{K}\left(\mathbf{z}_{i},\mathbf{z}^{\prime}_{j}\right)$}. The Gaussian noise on $\mathbf{y}\in\mathbb{R}^{1\times k}$ is typically handled by regularizing the Gram matrix with the noise variance $\sigma_{w}^{2}$, while the predictions of $d$ are noise free. Propagating the GP (4) over multiple steps is analytically intractable due to the non-Gaussian state distribution. Therefore, we use the typical approximations described in [hewing_cautious_2020, baltussen2025]. In addition, we employ a sparse GP [snelson_sparse_2005] with $M$ inducing points equally spaced over the MPC horizon.
@@ IV-A Gaussian Process Model Predictive Control
We consider a passive GP-MPC [baltussen2025], that we afterward extend with active learning. This passive GP-MPC uses the posterior mean and covariance to anticipate the future hyperstate. The passive dual GP-MPC is formulated as follows:
@@
 $\hskip-7.96677pt\text{ s.t. }\hat{x}_{i+1\mid k}$ $=f\left(\hat{x}_{i\mid k},\hat{u}_{i\mid k}\right)+Bd\left(\mathbf{z}_{i\mid k},\mathcal{I}_{k}\right),$
@@
Here, $\Sigma_{k}$ denotes the joint covariance of the state $\Sigma^{x}_{k}$ and GP $\Sigma^{d}_{k}$; for further details we refer the reader to [baltussen2025]. We use the state covariance to heuristically tighten the constraints to $\hat{\mathbb{X}}_{i\mid k}\bigl(\hat{x}_{i\mid k},\Sigma^{x}_{i\mid k}\bigr):=\{\hat{x}_{i\mid k}\in\mathbb{X}\ominus\mathbb{S}(2\sigma_{x,i\mid k})\}$, where $\sigma_{x,i\mid k}\in\mathbb{R}^{n_{x}}$ is the predicted standard deviation of the state and $\mathbb{S}$ is a hyperrectangle. Note that this MPC only leverages the cautious property of dual control and does not actively explore the state space. The posterior mean (6b) and covariance (6c) of the state define the causally anticipated hyperstate: $\xi_{i\mid i,k}:=\mathcal{N}(\hat{x}_{i\mid k},\Sigma^{x}_{i\mid k}),\quad i=1,2,\ldots,N,$ where $\xi_{i\mid i,k}$ depends on the causal anticipation of the information that will be available up to time $k+i$ computed at time $k\in\mathbb{N}$ based on the information $\mathcal{I}_{k}$, see Section III-A. As in [baltussen2025], we use a custom GP implementation in CasADi [andersson_casadi_2019] that preserves the sensitivities of the GP covariance. We use CasADi’s symbolic framework to compute the hyperstate gradients. Consequently, the GP-MPC conditions the covariance, and, hence, the hyperstate, on the control sequence $\hat{U}_{k}$, and can be solved using standard optimization solvers like IPOPT [wachter_implementation_2006]. The resulting policy features the dual effect.
@@
$\vskip-2.84526pt\Sigma^{x}_{i\mid k}:=\mathbb{E}\bigl[\Sigma_{k+i}^{x}\mid\mathcal{I}_{k},\hat{U}^{*}_{0:i-1\mid k}\bigr]\neq\mathbb{E}\bigl[\Sigma_{k+i}^{x}\mid\mathcal{I}_{k}\bigr].$

### v2

@@ Abstract
This manuscript presents a dual model predictive controller (MPC) that balances the two objectives of dual control, namely, system identification and control. In particular, we propose a Gaussian process (GP)-based MPC that uses the posterior GP covariance for active learning. The dual MPC can steer the system towards states with high covariance, or to the setpoint, thereby balancing system identification and control performance (exploration vs. exploitation). We establish robust constraint satisfaction of the novel dual MPC through a contingency plan. We demonstrate the dual MPC in a numerical study of a nonlinear system with nonparametric uncertainties.
@@ I Introduction
One of the reasons for the high interest in model predictive control (MPC), in particular for safety-critical systems, is its ability to systematically handle constraints on the states and inputs. Yet, the performance and safety of MPC heavily depend on the accuracy of the model of the system. In various applications, the system dynamics may be uncertain or partially unknown, and need to be identified online. This requires sufficiently informative data, and hence, the excitation of the system. However, these system excitations may conflict with the control objective and safety requirements. Balancing these objectives, identification and control, is the focus of dual control [FELDBAUM1963541]. Therefore, studying dual control in the context of MPC is of high relevance for simultaneous identification and control of safety-critical systems.
@@ I-A The Dual Effect
When controlling a system that is subject to uncertainties, a controller can only make decisions based on the available information from the system. However, the possibility of causally anticipating future measurements can be leveraged to improve state estimation or prediction, and hence, control performance. When a controller incorporates a ‘future measurement program’ and its associated statistics, this controller can causally anticipate these future measurements [bar-shalom_dual_1974]. The controller is said to have the dual effect, when, in addition to influencing the system state, it can also influence the state uncertainty, e.g., the state covariance. As a consequence, the dual effect is comprised of two properties [bar-shalom_dual_1974]: 1) Caution: the dependence of the state covariance on the control policy naturally introduces a degree of inherent caution – or a lack thereof. In the context of MPC, this dependence can be leveraged to tighten or relax the state constraints based on the anticipated state covariance. 2) Exploration: this anticipated state covariance can also be used to introduce an explorative or exciting element to the control policy. Again, in the context of MPC, the state covariance can be used in the MPC cost function to steer towards regions of high uncertainty that may be of interest for system identification and control performance. For a given control problem, the solution to the Bellman equation defines the globally optimal control policy, which naturally balances identification and control in an optimal way. In practice, however, solving the Bellman equation is generally intractable and requires approximate solutions [mesbah_stochastic_2018].
@@ I-B1 Dual Control
In implicit dual control, the MPC cost function is used to approximate the Bellman equation. These types of methods are typically based on and related to approximate dynamic programming [mesbah_stochastic_2018]. For example, [arcari_dual_2020a] and [HUSHARP2024] use scenario trees in MPC to address the trade-off between exploration and exploitation. This exploration and exploitation trade-off is then implicitly governed by the cost function [mesbah_stochastic_2018]. However, existing implicit dual MPC methods rely on additional safety filters [HUSHARP2024], are limited to linear systems [soloperto_dual_2019], or do not consider state constraints [arcari_dual_2020a], which is essential for safety-critical systems. The implementation and computational complexity of implicit dual MPC limit its practical applicability [mesbah_stochastic_2018].
In explicit dual control, the explorative element of the control policy is introduced explicitly, e.g., by enforcing a persistently exciting input signal, or adding a term to the cost function that minimizes (future) model uncertainty [mesbah_stochastic_2018]. We refer to [mesbah_stochastic_2018] for a detailed review of these methods. Although the exploration in explicit dual MPC typically requires some form of expert knowledge and is often heuristic [mesbah_stochastic_2018], these methods can explicitly account for state constraints. Therefore, explicit dual MPC is of high interest, especially for safety-critical systems. While existing methods are highly relevant, they mostly focus on systems with parametric uncertainties [mesbah_stochastic_2018] or fixed model structures [heirung_stochastic_2017]. This raises the question of how to effectively extend these methods to address nonparametric model uncertainties in dual MPC.
@@ I-B2 Gaussian Process MPC
Gaussian process (GP) regression is a Bayesian regression method that enables flexible, nonparametric modeling along with explicit uncertainty quantification, which can be leveraged with limited expert knowledge and few data points [kocijan_modelling_2016]. It has been successfully used to identify and control systems with unknown dynamics using MPC [hewing_cautious_2020, baltussen2025, dubied2025robustadaptivempcformulation]. While GP-MPC is often applied as an effective heuristic method [hewing_cautious_2020, baltussen2025], it can be complemented with robustly safe policies [geurts_baltussen_2025] that ensure constraint satisfaction under uncertainty. Alternatively, probabilisitic safety guarantees on GP-MPC are provided in [dubied2025robustadaptivempcformulation]. However, existing GP-MPC methods generally lack the dual effect as the GP covariance is not conditioned on the predicted control sequence. Typically, the GP covariance is fixed prior to solving the MPC [hewing_cautious_2020] or decoupled from the MPC in a zero-order optimization algorithm [dubied2025robustadaptivempcformulation, lahr2024l4acados] to enable the real-time deployment of GP-MPC. Although these works are of great interest, decoupling the GP covariance excludes the dual effect. To bridge this gap, [baltussen2025] presents a dual MPC that does condition the GP covariance on the predicted control sequence, yielding a controller with the dual effect, which is a neccessary property for active learning with GP-MPC.
@@ I-C Contributions
As mentioned before, in [baltussen2025] we introduced a GP-MPC that features the dual effect. However, we did not explicitly incentivize exploration. Hence, this passively learning dual MPC was limited to the cautious property of the dual effect. In this work, we extend the GP-MPC to actively explore the state space and to learn unknown system dynamics while ensuring robust constraint satisfaction. The main contributions of this manuscript are as follows. 1) We propose an explicit dual MPC that actively learns system dynamics using the active learning framework from [soloperto_augmenting_2020]. 2) We prove recursive feasibility of the actively learning MPC by incorporating a contingency horizon[alsterda_contingency_2019]. 3) We demonstrate the effectiveness of the proposed dual MPC through a numerical study on a nonlinear system with nonparametric, unknown dynamics. In addition, we compare our dual MPC with a robust MPC (RMPC), a passively learning dual MPC [geurts_baltussen_2025], and the active learning framework from [soloperto_augmenting_2020] with a single horizon.
@@ III-A Control Policies and the Dual Effect
According to Definition 2, only the $N$-measurement feedback and closed-loop control policies exhibit this dual effect [bar-shalom_dual_1974, mesbah_stochastic_2018]. This effect is a necessary property for active learning in GP-MPC, as it allows the MPC to actively influence the posterior GP covariance and steer the system toward uncertain regions to learn the dynamics and enhance performance.
@@ III-B Robust Model Predictive Control
$\bar{u}_{i\mid k}$ $\in\bar{\mathbb{U}}_{i},\qquad\qquad\hskip 5.69054pti=0,1,\dots,N-1,$ (3d)
$\bar{x}_{i\mid k}$ $\in\bar{\mathbb{X}}_{i},\qquad\qquad\hskip 5.69054pti=0,1,\dots,N,$ (3e)
@@ III-C1 Gaussian Process Regression
$y_{k+1}=B^{{\dagger}}\left(x_{k+1}-f(x_{k},u_{k})\right)=g\left(x_{k},u_{k}\right)+\tilde{v}_{k}$
with $B^{{\dagger}}=\left(B^{\top}B\right)^{-1}B^{\top}$ and $B^{{\dagger}}v_{k}=\tilde{v}_{k}\sim\mathcal{N}(0,\Sigma_{v})$. We use the training data $\mathbf{Z}_{k}=[\mathbf{z}_{0},\dots,\mathbf{z}_{k-1}]\in\mathbb{R}^{n_{z}\times k}$ and $\mathbf{y}_{k}=[y_{1},\dots,y_{k}]\in\mathbb{R}^{n_{g}\times k}$} at time $k\in\mathbb{N}_{\geq 1}$ in order to approximate the residual dynamics by a GP $\tilde{\mathbf{d}}\hskip-1.42262pt:\mathbb{R}^{n_{z}}\rightarrow\mathbb{R}^{n_{g}}$ such that $g\left(x_{k},u_{k}\right)\approx\tilde{\mathbf{d}}\left(\mathbf{z}_{k}\right)$. We confine ourselves to scalar GPs $(n_{g}=1)$ for the sake of readability. Secondly, we impose a prior distribution on the GP through a universal kernel function $\mathfrak{K}\left(\mathbf{z},\mathbf{z}^{\prime}\right)$ [kocijan_modelling_2016]. As we use the GP for model augmentation, we employ a zero mean prior. Conditioning the prior GP on the information $\mathcal{I}_{k}$ yields the posterior GP. Finally, we predict the residual dynamics over the MPC horizon by querying the posterior GP $\tilde{d}(\mathbf{z})\in\mathbb{R}$ at a predicted test point $\mathbf{z}_{i\mid k}\in\mathbb{R}^{n_{z}}$ for $i=0,1,\dots,N-1$ at time $k$ with
@@
$d(\mathbf{z})$ $=\mathbf{k}_{\mathbf{zZ}}\left(K_{\mathbf{ZZ}}+I\sigma^{2}_{v}\right)^{-1}\mathbf{y},$ (5a)
$\Sigma^{d}(\mathbf{z})$ $=\mathfrak{K}_{\mathbf{zz}}-\mathbf{k}_{\mathbf{zZ}}\left(K_{\mathbf{ZZ}}+I\sigma^{2}_{v}\right)^{-1}\mathbf{k}_{\mathbf{Z}\mathbf{z}}.$ (5b)
Here $\mathfrak{K}_{\mathbf{zz}^{\prime}}=\mathfrak{K}\left(\mathbf{z},\mathbf{z}^{\prime}\right)$, $\mathbf{k}_{\mathbf{Zz}}\in\mathbb{R}^{{k}}$ is the concatenation of the kernel function evaluated at the test point $\mathbf{z}$ and the training set $\mathbf{Z}$, where $\mathbf{k}_{\mathbf{Zz}}]_{i}=\mathfrak{K}\left(\mathbf{z}_{i},\mathbf{z}\right)$ and $\mathbf{k}_{\mathbf{Zz}}^{\top}=\mathbf{k}_{\mathbf{zZ}}$}, and $K_{\mathbf{ZZ}}\in\mathbb{R}^{{k}\times{k}}$ is a Gram matrix, which satisfies $K_{\mathbf{ZZ^{\prime}}}]_{ij}=\mathfrak{K}\left(\mathbf{z}_{i},\mathbf{z}^{\prime}_{j}\right)$}. The Gaussian noise on $\mathbf{y}\in\mathbb{R}^{1\times k}$ is typically handled by regularizing the Gram matrix with the noise variance $\sigma_{v}^{2}$, while the predictions of $d$ are noise free. Propagating the GP (4) over multiple steps is intractable as the nonlinear dynamics yield a non-Gaussian state distribution. Therefore, we use typical approximations to preserve the Gaussian distribution of the state, as described in [hewing_cautious_2020, baltussen2025]. In addition, we employ a sparse GP [snelson_sparse_2005] with $M$ inducing points equally spaced over the MPC horizon.
@@ IV-A Gaussian Process Model Predictive Control
We consider a passively learning GP-MPC [baltussen2025], that we afterward extend with active learning. This passive GP-MPC uses the posterior GP mean and covariance to anticipate the future hyperstate and adapt the state constraints accordingly. The passive dual GP-MPC is formulated as follows:
@@
 $\hskip-7.96677pt\text{ s.t. }\hat{x}_{i+1\mid k}$ $=f\left(\hat{x}_{i\mid k},\hat{u}_{i\mid k}\right)+Bd\left(\hat{\mathbf{z}}_{i\mid k},\mathcal{I}_{k}\right),$
@@
Here, $\Sigma_{i\mid k}$ denotes the joint covariance of the state $\Sigma^{x}_{i\mid k}$ and the GP $\Sigma^{d}_{i\mid k}$. For details we refer the reader to [baltussen2025]. We use the state covariance to heuristically adapt the constraints to $\hat{\mathbb{X}}_{i\mid k}(\hat{x}_{i\mid k},\Sigma^{x}_{i\mid k}):=\{\hat{x}_{i\mid k}\in\mathbb{X}\ominus\mathbb{S}(2\sigma_{x,i\mid k})\}$, where $\sigma_{x,i\mid k}\in\mathbb{R}^{n_{x}}$ is the predicted standard deviation of the state and $\mathbb{S}$ is a hyperrectangle. Note that this MPC only leverages the cautious property of dual control and does not actively explore the state space. The posterior mean (6b) and covariance (6c) of the state define the causally anticipated hyperstate: $\xi_{i\mid i,k}:=\mathcal{N}(\hat{x}_{i\mid k},\Sigma^{x}_{i\mid k}),i=1,2,\ldots,N,$ where $\xi_{i\mid i,k}$ depends on the causal anticipation of the information that will be available up to time $k+i$ computed at time $k\in\mathbb{N}$ based on the information $\mathcal{I}_{k}$, see Section III-A. As in [baltussen2025], we use a custom GP implementation in CasADi [andersson_casadi_2019] that preserves the sensitivities of the GP covariance. We use CasADi’s symbolic framework to compute the hyperstate gradients. Consequently, the GP-MPC conditions the covariance, and, hence, the hyperstate, on the control sequence $\hat{U}_{k}$, and can be solved using standard optimization solvers like IPOPT [wachter_implementation_2006]. The resulting policy features the dual effect.
@@
$\Sigma^{x}_{i\mid k}:=\mathbb{E}\bigl[\Sigma_{k+i}^{x}\mid\mathcal{I}_{k},\hat{U}^{*}_{0:i-1\mid k}\bigr]\neq\mathbb{E}\bigl[\Sigma_{k+i}^{x}\mid\mathcal{I}_{k}\bigr].$